14) sara bought 5000 stocks at a price of 2 USD and her broker asked her to deposit 6000 USD. sara then sold the stock at 2.5 USD . what is sara's rate of return
Answer #1

Dollar Return = (2.5 - 2)(5,000) = $2,500

Return Rate = 2,500/6,000

Return Rate = 41.67%